OpenStudy (anonymous):

A number minus 10 times its reciprocal is 3. Find the number.

OpenStudy (anonymous):

x-(10/x)=3

OpenStudy (anonymous):

x^2-10-3x=0 (x-5)(x+2)=0 x=5 or x=-2